To unload a truck a down ramp of length 8.0 m is arranged so it is downwards at an angle of
30.0° to the horizontal. Boxes of fruit and placed on the ramp and they slide down attaining a
velocity of 4.00 m/s by the time it reaches the bottom of the ramp. Determine the energy loss
due to friction while sliding down the ramp by considering the changes in the potential energy
and the kinetic energy of the box, between the top and bottom of the ramp. Hence determine a
value for the kinetic coefficient of friction
1
Expert's answer
2020-04-08T10:36:25-0400
"E_{lost}=mgh-0.5mv^2"

"E_{lost}=m(gl\sin{30}-0.5(v)^2)"

"E_{lost}=4((9.8)(8)(0.5)-0.5(4)^2)=125\ J"

"\mu=\frac{125}{4((9.8)(8)(0.5))}=0.80"